OpenStudy (anonymous):

Help! Graph y=tan(x+π/6)

OpenStudy (jdoe0001):

\(\large { \textit{function transformations} \\ \quad \\ \begin{array}{llll} \begin{array}{llll} shrink\ or\\ expand\\ by\ {\color{purple}{ A}}\cdot {\color{blue}{ B}}\end{array} \qquad \begin{array}{llll} vertical\\ shift\\ by \ {\color{green}{ D}} \end{array} \begin{array}{llll}{\color{green}{ D}} > 0& Up\\ {\color{green}{ D}} < 0 & Down\end{array} \\ \qquad \downarrow\qquad\qquad\quad \downarrow\\ y = {\color{purple}{ A}} ( {\color{blue}{ B}}x + {\color{red}{ C}} ) + {\color{green}{ D}}\\ \qquad\qquad\uparrow \\ \qquad\begin{array}{llll} horizontal\\ shift\\ by \ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\end{array} \begin{array}{llll}\frac{{\color{red}{ C}}}{{\color{blue}{ B}}} > 0 & to \ left\\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}} < 0& to \ right\end{array} \end{array} }\) notice, is just tan(x) horizontally shifted

OpenStudy (anonymous):

it's shifted to the left by π/6 right?

OpenStudy (jdoe0001):

yes